### v1

# Spreading in claw-free cubic graphs

Abstract
Let $p\in\mathbb{N}$ and $q\in\mathbb{N}\cup\{\infty\}$. We study a dynamic coloring of the vertices of a graph $G$ that starts with an initial subset $S$ of blue vertices, with all remaining vertices colored white. If a white vertex $v$ has at least $p$ blue neighbors and at least one of these blue neighbors of $v$ has at most $q$ white neighbors, then by the spreading color change rule the vertex $v$ is recolored blue. The initial set $S$ of blue vertices is a $(p,q)$-spreading set for $G$ if by repeatedly applying the spreading color change rule all the vertices of $G$ are eventually colored blue. The $(p,q)$-spreading set is a generalization of the well-studied concepts of $k$-forcing and $r$-percolating sets in graphs. For $q\geq 2$, a $(1,q)$-spreading set is exactly a $q$-forcing set, and the $(1,1)$-spreading set is a $1$-forcing set (also called a zero forcing set), while for $q=\infty$, a $(p,\infty)$-spreading set is exactly a $p$-percolating set. The $(p,q)$-spreading number, ${\sigma_{(p,q)}}(G)$, of $G$ is the minimum cardinality of a $(p,q)$-spreading set. In this paper, we study $(p,q)$-spreading in claw-free cubic graphs. While the zero-forcing number of claw-free cubic graphs was studied earlier, for each pair of values $p$ and $q$ that are not both $1$ we either determine the $(p,q)$-spreading number of a claw-free cubic graph $G$ or show that $\sigma_{(p,q)}(G)$ attains one of two possible values.
Keywords: Bootstrap percolation; zero forcing set; $k$-forcing set; spreading
AMS subject classification: 05C35, 05C75
1 Introduction
In this paper we continue the study of dynamic graph colorings and explore the concept of $(p,q)$-spreading in claw-free cubic graph. The concept of spreading in graphs is a generalization of the well-studied concepts of $k$-forcing and $r$-percolating sets in graphs.
Consider a dynamic coloring of the vertices of a graph $G$ that starts with an initial subset $S$ of blue vertices, with all remaining vertices colored white. For $k\in\mathbb{N}$, the color change rule in $k$-forcing is defined as follows: if a blue vertex $u$ has at most $k$ white neighbors, then all white neighbors of $u$ are recolored blue. In particular, when $k=1$, this is the color change rule for zero forcing in graphs. The initial set $S$ of blue vertices is a $k$-forcing set for $G$ if by repeatedly applying the color change rule in $k$-forcing all the vertices of $G$ are eventually colored blue. The $k$-forcing number of $G$, denoted $F_{k}(G)$, is the minimum cardinality among all $k$-forcing set of $G$. When $k=1$, the $k$-forcing number of $G$ is called the zero forcing number and is denoted by $Z(G)$, and so $Z(G)=F_{1}(G)$.
The concept of bootstrap percolation has a similar definition to that of zero forcing, yet a different motivation. It was introduced as a simplified model of a magnetic system in 1979 [8], and was later studied on random graphs and also on deterministic graphs. One of the first graphs in which bootstrap percolation was studied were grids [5, 6].
As before, we consider a dynamic coloring of the vertices of a graph $G$ that starts with an initial subset $S$ of blue vertices, with all remaining vertices colored white. We refer to blue vertices as “infected” and white vertices as “uninfected”. For $r\in\mathbb{N}$, the color change rule in $r$-percolation is defined as follows: if a white (uninfected) vertex $u$ has at least $r$ blue (infected) neighbors, then the vertex $u$ is recolored blue. The initial set $S$ of blue vertices is an $r$-neighbor bootstrap percolating set, or simply an $r$-percolating set, of $G$ if by repeatedly applying the color change rule in $r$-percolation all the vertices of $G$ are eventually colored blue. The $r$-neighbor bootstrap percolation number, or simply the $r$-percolation number, of $G$, denoted $m(G,r)$, is the minimum cardinality among all $r$-neighbor bootstrap percolating sets of $G$.
Motivated by the similarity of the definitions of the above two concepts, a common generalization of $k$-forcing and $r$-bootstrap percolation was introduced in [7]. Let $p\in\mathbb{N}$ and $q\in\mathbb{N}\cup\{\infty\}$. As before, we consider a dynamic coloring of the vertices of a graph $G$ that starts with an initial subset $S$ of blue vertices, with all remaining vertices colored white. The color change rule in $(p,q)$-spreading is defined as follows: if a white vertex $w$ has at least $p$ blue neighbors, and one of the blue neighbors of $w$ has at most $q$ white neighbors, then the vertex $w$ is recolored blue. The initial set $S$ of blue vertices is a $(p,q)$-spreading set of $G$, if by repeatedly applying the $(p,q)$-spreading color change rule all the vertices of $G$ are eventually colored blue. The $(p,q)$-spreading number, denoted ${\sigma_{(p,q)}}(G)$, of a graph $G$ is the minimum cardinality among all $(p,q)$-spreading sets.
The $(p,q)$-spreading set is a generalization of the well-studied concepts of $k$-forcing and $r$-percolating sets in graphs. For $q\geq 2$, a $(1,q)$-spreading set is exactly a $q$-forcing set [2], and the $(1,1)$-spreading set is a $1$-forcing set (also called a zero forcing set [1, 14]), while for $q=\infty$ and for any $q$ that is not smaller than the maximum degree of $G$, a $(p,q)$-spreading set is exactly a $p$-percolating set. In the seminal paper [7], the complexity of the decision version of the $(p,q)$-spreading number was proved to be NP-complete for all $p$ and $q$, while efficient algorithms for determining these numbers were found in trees. In addition, for almost all values of $p$ and $q$, the $(p,q)$-spreading numbers of Cartesian grids $P_{n}\mathbin{\Box}P_{m}$ were established in [7].
A graph is claw-free if it does not contain $K_{1,3}$ as an induced subgraph. A cubic graph (also called a $3$-regular graph) is a graph in which every vertex has degree $3$. In this paper we study spreading in claw-free cubic graphs.
1.1 Graph theory notation and terminology
For notation and graph theory terminology, we in general follow [11]. Specifically, let $G$ be a graph with vertex set $V(G)$ and edge set $E(G)$, and of order $n(G)=|V(G)|$ and size $m(G)=|E(G)|$. A neighbor of a vertex $v$ in $G$ is a vertex $u$ that is adjacent to $v$, that is, $uv\in E(G)$. The open neighborhood $N_{G}(v)$ of a vertex $v$ in $G$ is the set of neighbors of $v$, while the closed neighborhood of $v$ is the set $N_{G}[v]=\{v\}\cup N(v)$. We denote the degree of $v$ in $G$ by $\deg_{G}(v)=|N_{G}(v)|$. For a set $S\subseteq V(G)$, its open neighborhood is the set $N_{G}(S)=\cup_{v\in S}N_{G}(v)$, and its closed neighborhood is the set $N_{G}[S]=N_{G}(S)\cup S$.
For a set $S\subseteq V(G)$, the subgraph induced by $S$ is denoted by $G[S]$. Further, the subgraph of $G$ obtained from $G$ by deleting all vertices in $S$ and all edges incident with vertices in $S$ is denoted by $G-S$; that is, $G-S=G[V(G)\setminus S]$. If $S=\{v\}$, then we also denote $G-S$ simply by $G-v$. If $F$ is a graph, then an $F$-component of $G$ is a component isomorphic to $F$. We denote by $\alpha(G)$ the independence number of $G$, and so $\alpha(G)$ is the maximum cardinality among all independent sets of $G$. A vertex cover of $G$ is a set $S$ of vertices such that every edge of $G$ is incident with at least one vertex in $S$. The vertex covering number $\beta(G)$ (also denoted $\tau(G)$ in the literature), equals the minimum cardinality among all vertex covers of $G$. By the well-known Gallai theorem, $\alpha(G)+\beta(G)=n(G)$ holds in any graph $G$.
We denote the path, cycle, and complete graph on $n$ vertices by $P_{n}$, $C_{n}$, and $K_{n}$, respectively, and we denote the complete bipartite graph with partite sets of cardinality $n$ and $m$ by $K_{n,m}$. A triangle in $G$ is a subgraph isomorphic to $K_{3}$, whereas a diamond in $G$ is an induced subgraph of $G$ isomorphic to $K_{4}$ with one edge missing, denoted by $K_{4}-e$. A graph is diamond-free if it does not contain $K_{4}-e$ as an induced subgraph. For $k\geq 1$ an integer, we use the standard notation $[k]=\{1,\ldots,k\}$.
1.2 Main results and organization of the paper
Davila and Henning studied zero forcing in claw-free cubic graphs [9]. They established two upper bounds, which are summarized in the following result (for the definition of diamond-necklace $N_{2}$, see Section 2):
Theorem 1
([9]) If $G\neq K_{4}$ is a connected, claw-free, cubic graph of order $n$, then the following properties hold:
  1. (a)
    $Z(G)\leq\alpha(G)+1$;
  2. (b)
    $Z(G)\leq\frac{n}{3}+1$, unless $G=N_{2}$ in which case $Z(G)=\frac{1}{3}(n+4)$.
It is proved in [9] that both upper bounds in Theorem 1 are sharp. However, it is suspected that the bounds may be improvable if the order is large enough. In addition, it is not known whether there exists a polynomial-time algorithm to determine the exact zero forcing number of a claw-free cubic graph; see [9]. Note that $Z(G)=\sigma_{(1,1)}(G)$, and the mentioned result (Theorem 1(a)) is placed in the $(1,1)$-entry of Table 1.
From properties of connected, claw-free, cubic graphs $G$ which we discuss in Section 2, if $G\neq K_{4}$ then there is a unique partition of the vertex set $V(G)$ into subsets each of which induces either a triangle or a diamond. The number of these subsets, called units, in $G$ is denoted by $u(G)$. Table 1 summarizes the values of $(p,q)$-spreading numbers in claw-free cubic graphs $G\neq K_{4}$ for all possible $p$ and $q$ in terms of the parameters $\alpha(G)$, $\beta(G)$, $u(G)$ and $n(G)$. Besides the $(1,1)$-entry giving the mentioned upper bound on $\sigma_{(1,1)}(G)$ all other values in Table 1 are obtained in this paper. It is easy to see that $\sigma_{(1,2)}(G)=2$ for any claw-free cubic graph $G$. Indeed, any adjacent pair of vertices in $G$ is a $(1,2)$-spreading set in $G$. The results $\sigma_{(1,q)}(G)=1$ for any $q\geq 3$ and $\sigma_{(p,q)}(G)=n(G)$ for any $p\geq 4$ and any $q\in\mathbb{N}\cup\{\infty\}$ are trivial consequences of definitions.
$p$ $q$ 1 2 $\geq 3$
1 $\leq\alpha(G)+1$ 2 1
2 $u(G)+1$ or $u(G)+2$ $u(G)$ or $u(G)+1$ $u(G)$ or $u(G)+1$
3 $\beta(G)$ or $\beta(G)+1$ $\beta(G)$ $\beta(G)$
$\geq 4$ $n(G)$ $n(G)$ $n(G)$
Table 1: $(p,q)$-spreading numbers of claw-free cubic graphs $G$, $G\neq K_{4}$.
The paper is organized as follows. In Section 3, we consider $(3,q)$-spreading numbers in claw-free cubic graphs; their values can be found in Line 3 of Table 1. (Note that $3$-percolation coincides with $(3,q)$-spreading when $q\geq\Delta(G)$.) In Section 4, we deal with $(2,q)$-spreading numbers of claw-free cubic graphs; see Line 2 in Table 1. In the case of $(2,q)$-spreading, where $q\geq 3$, $\sigma_{(2,q)}(G)=u(G)+1$ precisely when $G$ is a diamond necklace $N_{k}$, while in all other claw-free cubic graphs $\sigma_{(2,q)}(G)=u(G)$. Thus the $2$-percolation number of claw-free cubic graphs is fully determined. On the other hand, when $q\leq 2$ a claw-free cubic graph can attain two values: $\sigma_{(2,2)}(G)\in\{u(G),u(G)+1\}$, and $\sigma_{(2,1)}(G)\in\{u(G)+1,u(G)+2\}$, while characterizing the families that attain one of the two possible values remains open. We conclude the paper with some open problems that arise from this study.
2 Claw-free cubic graphs
The following property of connected, claw-free, cubic graphs is established in [13].
Lemma 2
([13]) If $G\neq K_{4}$ is a connected, claw-free, cubic graph of order $n$, then the vertex set $V(G)$ can be uniquely partitioned into sets each of which induces a triangle or a diamond in $G$.
By Lemma 2, the vertex set $V(G)$ of connected, claw-free, cubic graph $G\neq K_{4}$ can be uniquely partitioned into sets each of which induce a triangle or a diamond in $G$. Following the notation introduced in [13], we refer to such a partition as a triangle-diamond partition of $G$, abbreviated $\Delta$-D-partition. We call every triangle and diamond induced by a set in our $\Delta$-D-partition a unit of the partition. A unit that is a triangle is called a triangle-unit and a unit that is a diamond is called a diamond-unit. (We note that a triangle-unit is a triangle that does not belong to a diamond.) Two units in the $\Delta$-D-partition are adjacent if there is an edge joining a vertex in one unit to a vertex in the other unit. In what follows, we define several structures in claw-free, cubic graphs that we will need when proving our main results. Some of these definitions have already appeared in the literature, see, for example, [3, 4, 13].
For $k\geq 2$ an integer, let $N_{k}$ be the connected cubic graph constructed as follows. Take $k$ disjoint copies $D_{1},D_{2},\ldots,D_{k}$ of a diamond, where $V(D_{i})=\{a_{i},b_{i},c_{i},d_{i}\}$ and where $a_{i}b_{i}$ is the missing edge in $D_{i}$. Let $N_{k}$ be obtained from the disjoint union of these $k$ diamonds by adding the edges $\{a_{i}b_{i+1}\colon i\in[k-1]\}$ and adding the edge $a_{k}b_{1}$. We call $N_{k}$ a diamond-necklace with $k$ diamonds. Let ${\cal N}_{\rm cubic}=\{N_{k}\mid k\geq 2\}$. A diamond-necklace, $N_{4}$, with four diamonds is illustrated in Figure 1.
Figure 1: A diamond-necklace $N_{4}$
For $k\geq 1$ an integer, let $F_{2k}$ be the connected cubic graph constructed as follows. Take $2k$ disjoint copies $T_{1},T_{2},\ldots,T_{2k}$ of a triangle, where $V(T_{i})=\{x_{i},y_{i},z_{i}\}$ for $i\in[2k]$. Let
$\begin{array}[]{lcl}E_{a}&=&\{x_{2i-1}x_{2i}\colon i\in[k]\}\vspace{0.1cm}\\ E_{b}&=&\{y_{2i-1}y_{2i}\colon i\in[k]\}\vspace{0.1cm}\\ E_{c}&=&\{z_{2i}z_{2i+1}\colon i\in[k]\},\end{array}$
where addition is taken modulo $2k$ (and so, $z_{1}=z_{2k+1}$). Let $F_{2k}$ be obtained from the disjoint union of these $2k$ triangles by adding the edges $E_{a}\cup E_{b}\cup E_{c}$. The resulting graph $F_{2k}$ we call a triangle-necklace with $2k$ triangles. Let $\mathcal{T}_{\rm cubic}=\{F_{2k}\colon k\geq 1\}$. The triangle-necklace $F_{6}$ is shown in Figure 2.
Figure 2: The triangle-necklace $F_{6}$
For $k\geq 2$ an integer, take $2k$ disjoint copies $T_{1},T_{2},\ldots,T_{2k}$ of a triangle, where $V(T_{i})=\{x_{i},y_{i},z_{i}\}$ for $i\in[2k]$, and take $k$ disjoint copies $D_{1},D_{2},\ldots,D_{k}$ of a diamond, where $V(D_{j})=\{a_{j},b_{j},c_{j},d_{j}\}$ and where $a_{j}b_{j}$ is the missing edge in $D_{j}$ for $j\in[k]$. Let
$\begin{array}[]{lcl}E_{1}&=&\{x_{2i-1}a_{i}\colon i\in[k]\}\vspace{0.1cm}\\ E_{2}&=&\{x_{2i}b_{i}\colon i\in[k]\}\vspace{0.1cm}\\ E_{3}&=&\{y_{i}z_{i+1}\colon i\in[k-1]\}\cup\{y_{k}z_{1}\},\vspace{0.1cm}\\ E_{4}&=&\{y_{i}z_{i+1}\colon i\in[2k-1]\setminus[k]\}\cup\{y_{2k}z_{k+1}\}.\end{array}$
Let $H_{2k}$ be obtained from the disjoint union of these $2k$ triangles and $k$ diamonds by adding the edges $E_{1}\cup E_{2}\cup E_{3}\cup E_{4}$. The resulting graph $H_{2k}$ we call a triangle-diamond-necklace with $k$ diamonds. Let $\mathcal{H}_{\rm cubic}=\{H_{2k}\colon k\geq 2\}$. The triangle-diamond-necklace $H_{6}$ is shown in Figure 3.
Figure 3: The triangle-diamond-necklace $H_{6}$
3 $3$-Percolation in claw-free, cubic graphs
In this section, we study $3$-percolation in claw-free, cubic graphs. The following lemma is already known in the literature and follows readily from the definition of $r$-bootstrap percolation.
Lemma 3
([12]) For $r\geq 2$ if $H$ is a subgraph of a graph $G$ such that every vertex in $H$ has strictly less than $r$ neighbors in $G$ that belong to $V(G)\setminus V(H)$, then every $r$-percolating set of $G$ contains at least one vertex of $H$.
As a consequence of Lemma 3, we establish a relationship between the $r$-percolation number and the independence number of an $r$-regular graph. Recall that $\beta(G)$ is the cardinality of a maximum vertex cover of $G$.
Proposition 4
Let $G$ be an $r$-regular graph of order $n$, where $r\geq 2$. A set $S\subset V(G)$ is an $r$-percolating set in $G$ if and only if $S$ is a vertex cover in $G$. In particular, $m(G,r)=\beta(G)=n-\alpha(G)$.
Proof. Let $S$ be an $r$-percolating set of an $r$-regular graph $G$. Suppose that $G-S$ contains an edge $uv$. Let $H=G[\{u,v\}]$, and so $H=K_{2}$ is a subgraph of $G$ such that every vertex in $H$ has strictly less than $r$ neighbors in $G$ that belong to $V(G)\setminus V(H)$. By Lemma 3, we infer that every $r$-percolating set of $G$ contains at least one vertex of $H$, which is a contradiction with our assumption. Thus, $S$ is a vertex cover of $G$. Conversely, suppose that $S$ is a vertex cover of $G$, and let $I=V(G)\setminus S$. Since the set $I$ is an independent set of $G$, every vertex in $I$ has all its $r$ neighbors in the set $V(G)\setminus I$. The set $S$ is therefore an $r$-percolating set of $G$. Since $m(G,r)$ is the cardinality of a maximum $r$-percolating set, we infer that $m(G,r)=\beta(G)$. Consequently, by Gallai’s theorem we also have $m(G,r)=n-\alpha(G)$$\Box$
As a consequence of Brook’s Coloring Theorem, we have the following trivial lower bound on the independence number of a graph.
Theorem 5
(Brooks) If $G\neq K_{n}$ is a connected graph of order $n$ with maximum degree $\Delta\geq 3$, then $\alpha(G)\geq\frac{n}{\Delta}$.
Proof. As a consequence of Brook’s Coloring Theorem, if $G\neq K_{n}$ is a connected graph of order $n$ with maximum degree $\Delta\geq 3$, then the chromatic number of $G$ is at most $\Delta$, that is, $\chi(G)\leq\Delta$. Alternatively, viewing a $\Delta$-coloring of $G$ as a partitioning of its vertices into $\Delta$ independent sets, called color classes, we infer by the Pigeonhole Principle that $G$ contains an independent set of cardinality at least $n/\Delta$, implying that $\alpha(G)\geq n/\Delta$$\Box$
The following upper bound on the independence number of a claw-free graph is given by several authors.
Theorem 6
([16, 10]) If $G$ is a claw-free graph of order $n$ with minimum degree $\delta$, then
$\alpha(G)\leq\left(\frac{2}{\delta+2}\right)n.$
As a consequence of Theorems 5 and 6, we have the following bounds on the independence number of a claw-free graph.
Theorem 7
If $G\neq K_{n}$ is a connected, claw-free graph of order $n$ with minimum degree $\delta$ and maximum degree $\Delta\geq 3$, then
$\frac{n}{\Delta}\leq\alpha(G)\leq\left(\frac{2}{\delta+2}\right)n.$
As a consequence of Proposition 4 and Theorem 7, we have the following result.
Theorem 8
If $G\neq K_{4}$ is a connected, claw-free, cubic graph of order $n$, then the following properties hold.
  1. (a)
    $\frac{1}{3}n\leq\alpha(G)\leq\frac{2}{5}n$.
  2. (b)
    $\frac{3}{5}n\leq m(G,3)\leq\frac{2}{3}n$.
We show next that the bounds of Theorem 8 are tight (in the sense that they hold for connected graphs of arbitrarily large orders). Suppose that $G\in\mathcal{T}_{\rm cubic}$. Thus, $G$ is a triangle-necklace $F_{2k}$ for some $k\geq 1$, and so $G$ has order $n=6k$ and contains $2k$ vertex disjoint triangles. Moreover, every unit of the (unique) $\Delta$-D-partition of $G$ is a triangle-unit, implying that $\alpha(G)\leq\frac{1}{3}n$. By Theorem 8(a), $\alpha(G)\geq\frac{1}{3}n$. Consequently, $\alpha(G)=\frac{1}{3}n$. For example, the white vertices in the triangle-necklace $F_{6}$ of order $n=18$ shown in Figure 4 form an $\alpha$-set of $F_{6}$ (of cardinality $6=\frac{1}{3}n$). By Proposition 4, we infer that $m(G,3)=\frac{2}{3}n$, and the corresponding set of shaded vertices in Figure 5 is a $\beta$-set of $G$. This shows that the lower bound of Theorem 8(a) is tight, as is the upper bound of Theorem 8(b).
Figure 4: A $\beta$-set in a triangle-necklace $F_{6}$
Suppose next that $G\in\mathcal{H}_{\rm cubic}$. Thus, $G$ is a triangle-diamond-necklace $H_{2k}$ for some $k\geq 2$, and so $G$ has order $n=10k$. We can choose an independent set of $G$ to contain one vertex from every triangle-unit and two vertices from every diamond-unit, implying that $\alpha(G)\geq 4k=\frac{2}{5}n$. For example, the white vertices in the triangle-diamond-necklace $H_{6}$ of order $n=30$ shown in Figure 5 form an $\alpha$-set of $H_{6}$ (of cardinality $12=\frac{2}{5}n$). By Proposition 4, we infer that $m(G,3)=\frac{3}{5}n$, and the corresponding set of shaded vertices in Figure 5 is a $\beta$-set of $H_{6}$ This shows that the upper bound of Theorem 8(a) is tight, as is the lower bound of Theorem 8(b).
Figure 5: A $\beta$-set in a triangle-diamond-necklace $H_{6}$
Note that $m(G,3)$ in (claw-free) cubic graphs coincides with $\sigma_{(3,q)}(G)$ for all $q\geq 3$, establishing the $(3,1)$-entry of Table 1. Hence, by Proposition 4, we have $\sigma_{(3,q)}(G)=m(G,3)=\beta(G)=n(G)-\alpha(G)$ for all $q\geq 3$. The result can be extended to $\sigma_{(3,2)}(G)$, which yields the following result.
Proposition 9
If $G$ is a connected, claw-free cubic graph of order $n$, then $\sigma_{(3,q)}(G)=\beta(G)$ for every $q\geq 2$.
Proof. We have already established that $\sigma_{(3,q)}(G)=\beta(G)$ for every $q\geq 3$, so it remains to resolve the case $q=2$. Consider a $3$-percolating set $P$ of a connected, claw-free cubic graph $G$, and color all its vertices blue. Note that $P$ is a vertex cover of $G$, and that $V(G)\setminus P$ is a maximum independent set. Now, since an independent set in $G$ contains at most one vertex from each triangle, and so a vertex cover contains at least two vertices in each triangle, we readily infer that every vertex in $P$ has a neighbor in $P$. Therefore, every (blue) vertex in $P$ has at most two (white) neighbors in $V(G)\setminus P$. Thus, $P$ is a $(3,2)$-spreading set. Since this is true for every $3$-percolating set $P$ of $G$, we infer that $\sigma_{(3,2)}(G)\leq m(G,3)$. Conversely, every $(3,2)$-spreading set of $G$ is necessarily a vertex cover of $G$, and is therefore a $3$-percolating set of $G$, implying that $m(G,3)\geq\sigma_{(3,2)}(G)$. Consequently, $\sigma_{(3,2)}(G)=m(G,3)=\beta(G)$$\Box$
We remark that $\sigma_{(3,1)}(G)>\beta(G)$ if $G\in\mathcal{H}_{\rm cubic}$ is a triangle-diamond necklace. Indeed, in any vertex cover $P$ of $G$, every vertex in $P$ has a neighbor in $P$, and so it does not have at most one neighbor in $V(G)\setminus P$.
Recall that two units in the $\Delta$-D-partition are adjacent if there exists at least one edge joining a vertex in one unit to a vertex in the other unit. We say that a unit in the $\Delta$-D-partition of $G$ is infected if all vertices in the unit are infected.
Proposition 10
If $G$ is a connected, claw-free cubic graph, then $\sigma_{(3,1)}(G)\leq\beta(G)+1$, and this bound is sharp.
Proof. Let $G$ be a connected, claw-free cubic graph, and let $P$ be a $3$-percolating set of $G$ satisfying $|P|=m(G,3)=\beta(G)$. We note that $P$ is a vertex cover of $G$ and $V(G)\setminus P$ is a maximum independent set of $G$. Clearly, every vertex, which is not in $P$, has three neighbors in $P$, and so the first condition of the $(3,1)$-spreading rule is always satisfied with respect to $P$. In addition, note that, since $P$ is a vertex cover, every triangle of $G$ has at least two vertices in $P$. Moreover, we can choose the set $P$ so that $P$ contains exactly two vertices from every triangle in $G$, or, equivalently, so that the independent set $V(G)\setminus P$ contains a vertex from every triangle in $G$. Now, consider an arbitrary triangle $T$ in $G$, and let $P^{*}=P\cup\{v\}$, where $v$ is the unique vertex in $V(T)\setminus P$. Let the vertices of $P^{*}$ be initially infected, and note that (with respect to $P^{*}$) the triangle $T$ is completely infected.
Let $u$ be a vertex that is not yet infected and is adjacent to an infected triangle $T^{\prime}$ (with all three vertices of $T^{\prime}$ infected). In particular, we note that $u\notin P^{*}$. Suppose firstly that $u$ is adjacent to two vertices of $T^{\prime}$. Thus, the vertices in $V(T^{\prime})\cup\{u\}$ induce a diamond-unit. In this case, the vertex $u$ immediately becomes infected, since both of its neighbors in $T^{\prime}$ have no other uninfected neighbor. As a result, the (unique) triangle containing the vertex $u$ becomes an infected triangle, and so the diamond-unit containing $u$ is infected. Suppose next that $u$ is adjacent to exactly one vertex $w^{\prime}$ of $T^{\prime}$. Once again in this case, the vertex $u$ immediately becomes infected, since its neighbor $w^{\prime}$ in $T^{\prime}$ has no other uninfected neighbor. On the other hand, suppose that $u$ is infected and is adjacent to an infected triangle $T^{\prime}$ (with all three vertices of $T^{\prime}$ infected), and there exists an uninfected neighbor $v$ of $u$ that belongs to the same unit as $u$. Then $v$ becomes infected, because $u$ has only one uninfected neighbor, namely $v$.
In all of the above cases, the unit, which is adjacent to an infected unit, becomes infected. Since $G$ is connected, and initially there is a triangle infected (making the unit in which the triangle lies also infected), we infer that due to the above arguments all vertices in $G$ become infected. Thus, $\sigma_{(3,1)}(G)\leq|P^{*}|=|P|+1=\beta(G)+1$, thereby proving the desired upper bound. The case of a triangle-diamond necklace $G\in\mathcal{H}_{\rm cubic}$ shows that the bound is sharp. $\Box$
4 $2$-Percolation in claw-free, cubic graphs
In this section, we study $2$-percolation in claw-free, cubic graphs. By Lemma 3, we infer that $m(K_{4},2)\geq 2$. However any two vertices in $K_{4}$ form a $2$-percolating set of the graph, and so $m(K_{4},2)\leq 2$. Consequently, $m(K_{4},2)=2$. In what follows we therefore restrict our attention to $2$-percolation in connected, claw-free, cubic graphs $G$ different from $K_{4}$. Thus, by Lemma 2, such a graph $G$ has a $\Delta$-D-partition (in which every unit is a triangle-unit or a diamond-unit). Recall that $u(G)$ is defined as the number of units in this (unique) $\Delta$-D-partition. In what follows, if $D$ is a diamond-unit in the $\Delta$-D-partition of $G$, then a dominating vertex in $D$ is a vertex that is adjacent to the three other vertices in the diamond-unit $D$.
Proposition 11
If $G\in{\cal N}_{\rm cubic}$, then $m(G,2)=u(G)+1$.
Proof. Suppose that $G\in{\cal N}_{\rm cubic}$, and so $G$ is a diamond-necklace $N_{k}$ for some $k\geq 2$. Thus, $G$ has $u(G)=k$ units, each of which is a diamond-unit. Let $S$ be a $2$-percolating set of $G$ of minimum cardinality, and so $S$ is an $2$-percolating set of $G$ satisfying $|S|=m(G,2)$. By Lemma 3, we infer that the set $S$ contains at least one vertex from every diamond-unit of $G$. Moreover if $D$ is a diamond-unit of $G$ and the set $S$ contains exactly one vertex from $D$, then by Lemma 3 we infer that such a vertex of $S$ is a dominating vertex in $D$. In particular, since $S$ contains at least one vertex from every unit in the $\Delta$-D-partition of $G$, we note that $m(G,2)=|S|\geq u(G)=k$.
As observed earlier, $|S|\geq k$. Suppose that $|S|=k$. By our earlier observations, the set $S$ contains exactly one vertex from every diamond-unit, namely the vertex from each diamond-unit that dominates that unit. The resulting set $S$ is a packing in $G$; that is, $d_{G}(u,v)\geq 3$ for every two distinct vertices $u$ and $v$ that belong to $S$. Moreover in this case, letting $H=G-S$, we note that $H$ is a cycle $C_{3k}$ and every vertex in $H$ has exactly one neighbor that belongs to $V(G)\setminus V(H)=S$. Thus by Lemma 3, we infer that every $2$-percolating set of $G$ contains at least one vertex of $H$. However this contradicts our supposition that $S$ is a $2$-percolating set of $G$ that contains no vertex of $H$. Hence, $m(G,2)=|S|\geq k+1$.
To establish an upper bound on $m(G,2)$ in this case when $G=N_{k}$, if $S^{*}$ consists of a dominating vertex from $k-1$ diamond-units and two dominating vertices from the remaining diamond-unit, then $S^{*}$ is a $2$-percolating set of $G$ and $|S^{*}|=k+1$. (For example, if $G=N_{4}$ (here $k=4$), then such a set $S^{*}$ illustrated in Figure 6 by the shaded vertices is a $2$-percolating set of $G$ and $|S^{*}|=5$.) This implies that $m(G,2)\leq|S^{*}|=k+1$. Consequently, $m(G,2)=k+1$. Thus in this case when $G\in{\cal N}_{\rm cubic}$, we have shown that $m(G,2)=u(G)+1$$\Box$
Figure 6: A diamond-necklace $N_{4}$ and a $2$-percolating set of size $5$
If two distinct units are joined by at least two edges, then we say that these two units are double-bonded.
Lemma 12
If $G\neq K_{4}$ is a connected, claw-free, cubic graph and $G\notin{\cal N}_{\rm cubic}$, then there exists a triangle-unit $T^{\prime}$, such that the graph $G-T^{\prime}$ has at most two components.
Proof. Let $\mathcal{T}$ be the set of all triangle-units in $G$. Suppose to the contrary that for every $T\in\mathcal{T}$ the graph $G-T$ consists of three components. Let $T_{1}\in\mathcal{T}$ be a triangle-unit such that at least one of the three components of the graph $G-T_{1}$ does not contain a triangle-unit. Denote this component by $G_{1}$ and let $D_{1}$ be the diamond-unit in $G_{1}$ adjacent to $T_{1}$. Then in $\cal T$, $D_{1}$ is adjacent to exactly one diamond-unit $D_{2}$, which is in turn adjacent to another diamond-unit $D_{3}$ and so on, which yields a contradiction since $G$ is finite. Therefore $G_{1}$ contains a triangle-unit $T^{\prime}\notin\mathcal{T}$, which is the final contradiction. $\Box$
Theorem 13
If $G\neq K_{4}$ is a connected, claw-free, cubic graph of order $n$ that contains $u(G)$ units, then
$m(G,2)=\left\{\begin{array}[]{ll}u(G)+1,&\mbox{if $G\in{\cal N}_{\rm cubic}$};\vspace{0.1cm}\\ u(G),&\mbox{otherwise}.\end{array}\right.$
Proof. By Lemma 3, we infer that every $2$-percolating set of $G$ contains at least one vertex from every triangle-unit and every diamond-unit of $G$. Moreover if $D$ is a diamond-unit of $G$ and a $2$-percolating set contains exactly one vertex from $D$, then by Lemma 3 we infer that such a vertex is a dominating vertex in $D$. In particular, since every $2$-percolating set of $G$ contains at least one vertex from every unit in the $\Delta$-D-partition of $G$, we note that $m(G,2)\geq u(G)$. If $G\in{\cal N}_{\rm cubic}$, then by Proposition 11, $m(G,2)=u(G)+1$. Hence we may assume that $G\notin{\cal N}_{\rm cubic}$, for otherwise the desired result follows.
Since every triangle-unit contributes $3$ to the order and every diamond-unit contributes $4$ to the order, we observe that if $G$ has order $n$ with $u_{t}$ triangle-units and $u_{d}$ diamond-units, then $u(G)=u_{t}+u_{d}$ and $n=3u_{t}+4u_{d}$. By assumption, $u_{t}\geq 1$. Thus, since $n$ is even, $u_{t}\geq 2$, that is, $G$ contains at least two triangle-units. According to Lemma 12, there exists a triangle-unit $T_{1}$ of $G$, where $V(T_{1})=\{t_{1},t_{2},t_{3}\}$, such that $G-T_{1}$ consists of at most two components. This implies that at most one of the edges incident with exactly one vertex of $T_{1}$ is a bridge. Since $G$ is connected, the triangle-unit $T_{1}$ is adjacent to at least one other unit. Let $U_{1}$ be a unit that is adjacent to $T_{1}$ such that the edge between $U_{1}$ and $T_{1}$ is not a bridge. Finally denote as $G_{1}$ the component of $G-T_{1}$ containing $U_{1}$.
As observed earlier, $m(G,2)\geq u(G)$. Hence it suffices for us to show that $m(G,2)\leq u(G)$. For this purpose, we construct a $2$-percolating set $S$ of $G$ that contains exactly one vertex from each unit of $G$. Initially, we let $S=\emptyset$. We consider two cases.
Case 1. The unit $U_{1}$ is a triangle-unit. Let $V(U_{1})=\{a_{1},b_{1},c_{1}\}$ and where $t_{1}c_{1}$ is an edge. In this case, we add to $S$ the vertices $t_{1}$ and $a_{1}$, and so $S=\{t_{1},a_{1}\}$. The vertices in $S$ are indicated by the shaded vertices in Figure 7(a). Since the vertex $c_{1}$ is adjacent to two infected vertices, namely vertices $t_{1}$ and $a_{1}$, the vertex $c_{1}$ becomes infected in the $2$-percolation process. Thereafter, $b_{1}$ is adjacent to two infected vertices, namely vertices $a_{1}$ and $c_{1}$, and so the vertex $b_{1}$ becomes infected in the $2$-percolation process. Thus, every vertex in the triangle-unit $U_{1}$ becomes infected.
[…]
Figure 7: Possible adjacent units in the proof of Theorem 13
Case 2. The unit $U_{1}$ is a diamond-unit. Let $V(U_{1})=\{a_{1},b_{1},c_{1},d_{1}\}$ and where $a_{1}b_{1}$ is the missing edge in $U_{1}$ and where $t_{1}b_{1}$ is an edge of $G$. In this case, we add to $S$ the vertices $t_{1}$ and $c_{1}$, and so $S=\{t_{1},c_{1}\}$. The vertices in $S$ are indicated by the shaded vertices in Figure 7(b). Since the vertex $b_{1}$ is adjacent to two infected vertices, namely vertices $t_{1}$ and $c_{1}$, the vertex $b_{1}$ becomes infected in the $2$-percolation process. Thereafter, $d_{1}$ is adjacent to two infected vertices, namely vertices $b_{1}$ and $c_{1}$, and so the vertex $d_{1}$ becomes infected in the $2$-percolation process. Once $d_{1}$ becomes infected, the vertex $a_{1}$ is adjacent to two infected vertices, namely vertices $c_{1}$ and $d_{1}$, and so the vertex $a_{1}$ becomes infected in the $2$-percolation process. Thus, every vertex in the diamond-unit $U_{1}$ becomes infected.
Cases 1 and 2 occur in Step 1 of our procedure to construct the set $S$. Thus after Step 1, the initial set $S$ consists of two vertices, namely a vertex $t_{1}$ from the triangle-unit $T_{1}$ and one vertex from the unit $U_{1}$ that is adjacent to $T_{1}$. Moreover in both cases, every vertex in the unit $U_{1}$ becomes infected. We now proceed to Step 2 of our procedure to construct the set $S$.
Suppose that the units $T_{1}$ and $U_{1}$ are double bonded. Thus there is a vertex $u\in U_{1}$ that is joined to a vertex of $T_{1}$ different from $t_{1}$. Renaming the vertices $t_{2}$ and $t_{3}$ if necessary, we may assume that $ut_{2}$ is such an edge between the units $T_{1}$ and $U_{1}$. In particular, we note that $t_{2}\notin S$. Since the vertex $t_{2}$ is adjacent to two infected vertices, namely vertices $u$ and $t_{1}$, the vertex $t_{2}$ becomes infected in the $2$-percolation process. Thereafter, $t_{3}$ is adjacent to two infected vertices, namely vertices $t_{1}$ and $t_{2}$, and so the vertex $t_{3}$ becomes infected in the $2$-percolation process. Thus, every vertex in the unit $T_{1}$ becomes infected. In this case, the set $S$ consists of two vertices, one vertex from each of the units $U_{1}$ and $T_{1}$.
Suppose that the units $T_{1}$ and $U_{1}$ are not double bonded. Let $U_{2}$ be a unit different from $T_{1}$ that is adjacent to the infected unit $U_{1}$, and let $uv$ be an edge that joins a vertex $u\in U_{1}$ and a vertex $v\in U_{2}$. Since $U_{2}\neq T_{1}$, we note that $v\neq t_{1}$. In particular, we note that $v\notin S$ since at this stage of the construction the set $S$ only contains the vertex $t_{1}$ and one other vertex, namely a vertex from $U_{1}$.
Suppose that $U_{2}$ is a triangle-unit. Let $V(U_{2})=\{v,v_{1},v_{2}\}$. We now add to the set $S$ exactly one of the vertices $v_{1}$ and $v_{2}$. By symmetry, we may assume that $v_{1}$ is added to the set $S$. Since the vertex $v$ is adjacent to two infected vertices, namely vertices $u$ and $v_{1}$, the vertex $v$ becomes infected in the $2$-percolation process. Thereafter, $v_{2}$ is adjacent to two infected vertices, namely vertices $v$ and $v_{1}$, and so the vertex $v_{2}$ becomes infected in the $2$-percolation process. Thus, every vertex in the triangle-unit $U_{2}$ becomes infected.
Suppose next that $U_{2}$ is a diamond-unit. Let $V(U_{2})=\{v,v_{1},v_{2},v_{3}\}$, where $v_{1}$ and $v_{2}$ are the dominating vertices of $U_{2}$ (and so, $vv_{3}$ is the missing edge in $D_{v}$). We now add to the set $S$ exactly one of the dominating vertices of $U_{2}$. By symmetry, we may assume that dominating vertex $v_{1}$ of $U_{2}$ is added to $S$. Since the vertex $v$ is adjacent to two infected vertices, namely vertices $u$ and $v_{1}$, the vertex $v$ becomes infected in the $2$-percolation process. Thereafter, $v_{2}$ is adjacent to two infected vertices, namely vertices $v$ and $v_{1}$, and so the vertex $v_{2}$ becomes infected in the $2$-percolation process. The vertex $v_{3}$ is now adjacent to two infected vertices, namely vertices $v_{1}$ and $v_{2}$, and so the vertex $v_{3}$ becomes infected in the $2$-percolation process. Thus, every vertex in the diamond-unit $U_{2}$ becomes infected.
After Step 2 of our procedure to construct the set $S$, every vertex in the unit $U_{2}$ becomes infected, and the set $S$ contains exactly one vertex from each of the units $U_{1}$ and $U_{2}$. Moreover either $T_{1}=U_{2}$, in which case $|S|=2$ and $S$ contains exactly one vertex from each of $U_{1}$ and $U_{2}$, or $T_{1}\neq U_{2}$, in which case $|S|=3$ and $S$ contains one vertex from each of the units $U_{1}$, $U_{2}$ and $T_{1}$.
Suppose that $i\geq 2$ and after Step $i$ of our procedure to construct the set $S$, the units $U_{1},U_{2},\ldots,U_{i}$ are all infected, and the set $S$ contains exactly one vertex from each of the units $U_{1},U_{2},\ldots,U_{i}$. Moreover if $T_{1}=U_{j}$ for some $j\in[i]\setminus\{1\}$, then after Step $i$ we have $|S|=i$ and the set $S$ contains exactly one vertex from each of the units $U_{1},U_{2},\ldots,U_{i}$, while if $T_{1}\neq U_{j}$ for any $j\in[i]$, then $|S|=i+1$ and $S$ contains one vertex from each of the units $U_{1},U_{2},\ldots,U_{i},T_{1}$.
In Step $i+1$, let $U_{i+1}$ be a unit different the units $U_{1},U_{2},\ldots,U_{i}$ that is adjacent to an infected unit $U_{j}$ for some $j\in[i]$, and let $uv$ be an edge that joins a vertex $u\in U_{j}$ and a vertex $v\in U_{i+1}$.
Suppose that $U_{i+1}$ is a triangle-unit. Let $V(U_{i+1})=\{v,v_{1},v_{2}\}$. In this case, either $U_{i+1}=T_{1}$ and renaming vertices of $T_{1}$ if necessary we may assume that $v_{1}=t_{1}$, or $U_{i+1}\neq T_{1}$ and we add to the set $S$ the vertex $v_{1}$. Since the vertex $v$ is adjacent to two infected vertices, namely vertices $u$ and $v_{1}$, the vertex $v$ becomes infected in the $2$-percolation process. Thereafter, $v_{2}$ is adjacent to two infected vertices, namely vertices $v$ and $v_{1}$, and so the vertex $v_{2}$ becomes infected in the $2$-percolation process. Thus, every vertex in the triangle-unit $U_{i+1}$ becomes infected.
Suppose next that $U_{i+1}$ is a diamond-unit. Let $V(U_{i+1})=\{v,v_{1},v_{2},v_{3}\}$, where $v_{1}$ and $v_{2}$ are the dominating vertices of $U_{i+1}$ (and so, $vv_{3}$ is the missing edge in $D_{v}$). We now add to the set $S$ the dominating vertex $v_{1}$ of $U_{i+1}$. Since the vertex $v$ is adjacent to two infected vertices, namely vertices $u$ and $v_{1}$, the vertex $v$ becomes infected in the $2$-percolation process. Thereafter by our earlier observations, the vertices $v_{2}$ and $v_{3}$ become infected in turn. Thus, every vertex in the diamond-unit $U_{i+1}$ becomes infected.
Continuing in the described way, the above process continues until every unit of $G_{1}$ and also the unit $T_{1}$ become infected. If $G-T_{1}$ is connected, then all vertices of $G$ have become infected and $S$ contains exactly one vertex from each unit of $G$, thus the proof is complete. Otherwise, if $G-T_{1}$ is not connected, then it has at most two components, and let $G_{2}$ be the component of $G-T_{1}$ different from $G_{1}$. Let $U_{1}^{\prime}$ be the unit in $G_{2}$ that is adjacent to $T_{1}$, and let $v_{2}a$ be the bridge of $G$ connecting $T_{1}$ with $U_{1}^{\prime}$. Now, adding a dominating vertex $b$ of $U_{1}^{\prime}$, different from $a$, we infer that all vertices of $U_{1}^{\prime}$ become infected (using the analogous arguments as in the previous cases). Now, we can continue with the same process, continuing by infecting the units $U_{i}^{\prime}$ in $G_{2}$ that are adjacent to an already infected unit by adding to $S$ appropriately selected dominating vertex of $U_{i}^{\prime}$. Since $G_{2}$ is connected, we infer that all vertices of $G_{2}$ become infected, where we put in $S$ exactly one vertex of each unit. Thus, $S$ is a $2$-percolating set of $G$, implying that $m(G,2)\leq u(G)$. As observed earlier, $m(G,2)\geq u(G)$. Consequently, $m(G,2)=u(G)$. This completes the proof of Theorem 13$\Box$
By definition of a $(p,q)$-spreading set in a graph, if $G$ is cubic graph, then a set is a $(2,3)$-spreading set if and only if it is $2$-percolating set, implying that $\sigma_{(2,3)}(G)=m(G,2)$. As an immediate consequence of Theorem 13, we infer the following result.
Corollary 1
If $G\neq K_{4}$ is a connected, claw-free, cubic graph of order $n$ that contains $u(G)$ units, then
$\sigma_{(2,3)}(G)=\left\{\begin{array}[]{ll}u(G)+1,&\mbox{if $G\in{\cal N}_{\rm cubic}$};\vspace{0.1cm}\\ u(G),&\mbox{otherwise}.\end{array}\right.$
We are now in a position to establish the following relationships between the $(2,2)$-spreading number and the $(2,3)$-spreading number of a connected, claw-free, cubic graph.
Proposition 14
If $G=K_{4}$ or if $G\in{\cal N}_{\rm cubic}$, then $\sigma_{(2,2)}(G)=\sigma_{(2,3)}(G)$.
Proof. If $G=K_{4}$, then $\sigma_{(2,2)}(G)=\sigma_{(2,3)}(G)=2$. Hence we may assume that $G\neq K_{4}$. Suppose that $G\in{\cal N}_{\rm cubic}$. By Corollary 1, we have $\sigma_{(2,3)}(G)=u(G)+1$. Moreover, as shown in the proof of Proposition 11, if $G=N_{k}$ for some $k\geq 2$, then we can choose the $(2,3)$-spreading set $S$ of $G$ to consists of a dominating vertex from $k-1$ diamond-units and two dominating vertices from the remaining diamond-unit in $G$ (as illustrated in Figure 6 in the case when $k=4$). However such a set $S$ is also a $(2,2)$-spreading set $S$ of $G$, and so $\sigma_{(2,2)}(G)\leq|S|=\sigma_{(2,3)}(G)$. As observed earlier, $\sigma_{(2,3)}(G)\leq\sigma_{(2,2)}(G)$. Consequently, if $G\in{\cal N}_{\rm cubic}$, then $\sigma_{(2,2)}(G)=\sigma_{(2,3)}(G)$$\Box$
Proposition 15
If $G$ is a connected, claw-free, cubic graph, then $\sigma_{(2,2)}(G)\leq\sigma_{(2,3)}(G)+1$.
Proof. Let $G$ be a connected, claw-free, cubic graph. If $G=K_{4}$ or if $G\in{\cal N}_{\rm cubic}$, then by Proposition 14 we have $\sigma_{(2,2)}(G)=\sigma_{(2,3)}(G)$. Hence we may assume that $G\neq K_{4}$ and $G\notin{\cal N}_{\rm cubic}$. To prove that $\sigma_{(2,2)}(G)\leq\sigma_{(2,3)}(G)+1$, let $S$ be a $(2,3)$-spreading set of $G$ as constructed in the proof of Theorem 13. Adopting the notation in the proof of Theorem 13, in Step 1 of our procedure to construct the set $S$ we add one additional vertex to the set $S$ as follows. If the unit $U_{1}$ is a triangle-unit (see Figure 7(a)), then we add the common neighbor of $t_{1}$ and $a_{1}$, namely the vertex $c_{1}$, to the set $S$. If the unit $U_{1}$ is a diamond-unit (see Figure 7(b)), then we add the common neighbor of $t_{1}$ and $c_{1}$, namely the vertex $b_{1}$, to the set $S$. Let $S^{*}$ denote the resulting set upon completion of the construction of the set $S$, and so $|S^{*}|=|S|+1=\sigma_{(2,3)}(G)+1$. With the addition of the vertex $c_{1}$ in Case 1 or the vertex $b_{1}$ in Case 2, the set $S^{*}$ is a $(2,2)$-spreading set $S$ of $G$, implying that $\sigma_{(2,2)}(G)\leq|S^{*}|=|S|+1=\sigma_{(2,3)}(G)+1$$\Box$
Every $(2,3)$-spreading set of $G$ is by definition also a $(2,2)$-spreading set of $G$, and so $\sigma_{(2,3)}(G)\leq\sigma_{(2,2)}(G)$. Hence as a consequence of Proposition 15, we have the following result.
Corollary 2
If $G$ is a connected, claw-free, cubic graph, then
$\sigma_{(2,3)}(G)\leq\sigma_{(2,2)}(G)\leq\sigma_{(2,3)}(G)+1.$
In addition, $u(G)\leq\sigma_{(2,2)}(G)\leq u(G)+1$.
If $G\in\mathcal{T}_{\rm cubic}$ then $\sigma_{(2,2)}(G)=u(G)$, and if $G\in{\cal N}_{\rm cubic}$ then $\sigma_{(2,2)}(G)=u(G)+1$. In both of these cases it holds that $\sigma_{(2,2)}(G)=\sigma_{(2,3)}(G)$, however this does not hold in general. For instance, see the graph $G$ in Figure 8 for which we have $\sigma_{(2,3)}=u(G)$, where the shaded vertices indicated in the figure form a $(2,3)$-spreading set of $G$. Yet, $\sigma_{(2,2)}=u(G)+1$.
Figure 8: A graph $G$ with $\sigma{(2,3)}=u(G)$ and $\sigma{(2,2)}=u(G)+1$
Finally, we consider $(2,1)$-spreading, and bound the corresponding invariant in claw-free cubic graphs. Again we see that it can achieve only two possible values.
Theorem 16
If $G\neq K_{4}$ is a connected, claw-free, cubic graph of order $n$ that contains $u(G)$ units, then $u(G)+1\leq\sigma_{(2,1)}(G)\leq u(G)+2$.
Proof. First note that if $S$ is a set of vertices that contains exactly one vertex from each unit of $G\neq K_{4}$, then every vertex $v\in S$ has at least two neighbors in $V(G)\setminus S$, and therefore such a set $S$ cannot be a $(2,1)$-spreading set. Hence, $u(G)+1\leq\sigma_{(2,1)}(G)$.
We show next that $\sigma_{(2,1)}(G)\leq u(G)+2$ by constructing a $(2,1)$-spreading set $S$ of $G$ satisfying $|S|\leq u(G)+2$. Let $T$ be an arbitrary triangle in $G$. Initially, we let $S=V(T)$, and so the vertices in $T$ form the set of initially infected vertices. We now enlarge the set $S$ as follows. Suppose that $T$ belongs to a diamond-unit $D$. In this case, we let $v$ be the vertex in the diamond-unit $D$ that is not in $T$. The vertex $v$ immediately becomes infected in the $(2,1)$-spreading process, since both of its neighbors in $T$ have no other uninfected neighbor. As a result, all vertices in the diamond-unit $D$ become infected, that is, $D$ becomes an infected unit. Suppose next that $T$ is a triangle-unit of $G$. Then $T$ is already an infected unit.
Finally let $T^{\prime}$ be the infected unit containing $T$ (that is, $T^{\prime}=D$ or $T^{\prime}=T$). We now extend the set $S$ in the same manner as in the proof of Theorem 13. Let $U$ be a unit adjacent to $T^{\prime}$. If $U$ is a triangle-unit, then we add to $S$ a vertex in $U$ that is not adjacent to any vertex in the unit $T^{\prime}$. If $U$ is a diamond-unit, then we add to $S$ a vertex in $U$ that is a dominating vertex of the unit $U$. Proceeding analogously as in the proof of Theorem 13, this results in all vertices of the unit $U$ becoming infected in the $(2,1)$-spreading process.
Continuing in this way by considering a vertex not yet infected and is adjacent to an infected triangle, we obtain a $(2,1)$-spreading set $S$ of $G$ starting with the set $V(T)$ and adding exactly one vertex from every unit of $G$ that does not contain the triangle $T$. Thus, $\sigma_{(2,1)}(G)\leq|S|=u(G)+2$$\Box$
5 Concluding remarks
We end this paper with some remarks concerning the computational complexity of determining $\sigma_{(p,q)}(G)$ in claw-free cubic graphs $G$. We have already mentioned that it is unclear whether one can determine the zero forcing number of a claw-free cubic graph efficiently. For some of the values of $p$ and $q$, where $p>1$ or $q>1$, there exists a polynomial-time algorithm to determine $\sigma_{(p,q)}(G)$. First, determining the independence number in claw-free graphs can be done in polynomial time (see [15], where a polynomial-time algorithm is presented even for the weighted version of the problem). Therefore the problem of determining $\sigma_{(3,q)}(G)$ is polynomial in claw-free cubic graphs $G$ for all $q\geq 2$. Similarly, one can efficiently determine the number of units in a claw-free cubic graph. Thus the problem of determining $\sigma_{(2,q)}(G)$ is also polynomial in claw-free cubic graphs $G$ for any $q\geq 3$.
The following problems remain open:
Problem 1
Provide a structural characterization of the connected claw-free cubic graphs $G$ for which $\sigma_{(2,2)}=u(G)$ (resp., $u(G)+1$). Is there a polynomial-time algorithm to recognize the corresponding classes of connected claw-free cubic graphs?
Problem 2
Provide a structural characterization of the connected claw-free cubic graphs $G$ for which $\sigma_{(2,1)}=u(G)+1$ (resp., $u(G)+2$). Is there a polynomial-time algorithm to recognize the corresponding classes of connected claw-free cubic graphs?
Problem 3
Provide a structural characterization of the connected claw-free cubic graphs $G$ for which $\sigma_{(3,1)}=\beta(G)$ (resp., $\beta(G)+1$). Is there a polynomial-time algorithm to recognize the corresponding classes of connected claw-free cubic graphs?
Acknowledgements
Boštjan Brešar was supported by the Slovenian Research Agency (ARRS) under the grants P1-0297, N1-0285, and J1-4008. Jaka Hedžet was supported by the Slovenian Research Agency (ARRS) under the grant P1-0297. Michael A. Henning was supported in part by the South African National Research Foundation (grants 132588, 129265) and the University of Johannesburg.
Declarations
Conflict of interest The authors do not have any financial or non financial interests that are directly or indirectly related to the work submitted for publication.
Data availability No data was used for the research described in this paper.